# Wonderful Team: Zero-Shot Physical Task Planning with Visual LLMs

### v5

@@ Abstract
We introduce Wonderful Team, a multi-agent Vision Large Language Model (VLLM) framework for executing high level robotic planning in a zero-shot regime. In our context, zero-shot high-level planning means that for a novel environment, we provide a VLLM with an image of the robot’s surroundings and a task description, and the VLLM outputs the sequence of actions necessary for the robot to complete the task. Unlike previous methods for high-level visual planning for robotic manipulation, our method uses VLLMs for the entire planning process, enabling a more tightly integrated loop between perception, control, and planning. As a result, Wonderful Team’s performance on a real-world semantic and physical planning tasks often exceeds methods that rely on separate vision systems. For example, we see an average 40% success-rate improvement on VimaBench over prior methods such as NLaP, an average 30% improvement over Trajectory Generators on tasks from the Trajectory Generator paper including drawing and wiping a plate, and an average 70% improvement over Trajectory Generators on a new set of semantic reasoning tasks including environment re-arrangement with implicit linguistic constraints. We hope these results highlight the rapid improvements of VLLMs in the past year, and motivate the community to consider VLLMs as an option for some high-level robotic planning problems in the future.
@@ 1 Introduction
High-level robotic planning is a difficult task to pin down. It requires an extensive knowledge of the physical world. If planning is to be done via language instructions, it also requires a semantic understanding of the language itself, and how the objects in the environment relate to sometimes vague linguistic instructions. Consider, for example, the task “Pick up the object to the right of the grapes and put it in the box.” To solve this task, we need to identify ‘the object to the right of the grapes,’ which is a non-trivial visual relationship inference problem. We also need to consider the implicit constraints of the environment. For example, if the box has a lid and is currently closed, we need to infer that the lid must first be removed before proceeding with the task.
@@
Figure 1: Comparison of high-level planning approaches for a color-matching fruit placement task. MOKA’s planned trajectory (b) shows several limitations: its unverified high-level plan generates suboptimal subgoals due to the lack of self correction (step 2 treating grape as green) and shows misalignment between VLLM-generated targets and visual grounding. For instance, at step 3, it identifies the orange fruit to be the ‘orange area’, and at step 4, it fails to properly pick the green apple due to not specifying the color. We discuss more on comparision with MOKA in Appendix F.4.
@@
Recently, Large Language Models (LLMs) have emerged as an enticing pathway for achieving high-level robotic planning. LLMs already have a built-in world model, including priors over object names and relative relationships. They also have the ability to understand and parse semantic meaning from language instructions. Several promising pipelines exist for converting language commands to high-level robotics plans Kwon et al. (2024). Typically, if vision-based planning is required, these pipelines make use of an off-the-shelf vision encoder such as LangSAM or OWL-ViT that finds objects in an environment and delivers their coordinates in pixel space. While this is a great idea in theory, we find that, in practice, such reliance on an external vision system often results in pipelines that are brittle and unable to recover from mistakes originating in the visual system.
@@
In this paper, we consider the problem of using Visual Large Language Models (VLLMs) to generate action-level robotics plans directly from environmental images. Compared to prior works, we find having VLLMs as the backbone of both planning and grounding in the pipeline to have several advantages. First, VLLMs combine the semantic reasoning power of language models with robust visual understanding, bridging the traditional gap between high-level planning and visual grounding. Thus, they do not struggle with commands such as “pick the orange and place it in the orange area.” The model can read the environment image and understand implicitly to treat the orange square as the orange area instead of the orange fruit, although both can be considered “orange area” without context. More importantly, VLLMs enable self-correction at both the planning and perception levels. For planning, if the initial high-level plan overlooks environmental constraints, VLLMs can identify these errors and suggest corrective actions. For perception, while traditional vision models fail permanently when objects cannot be found, our VLLM-based perception module can refine its search strategy through reasoning about alternative viewpoints and object appearances. This ability to self-correct improves success rates on real-world spatial and semantic planning tasks by approximately 80% versus a baseline GPT-4o model with no self-correction, with many tasks going from a 0% success rate to a 100% success rate after self-correction is introduced.
Our contributions can be summarized as follows
@@
  • Zero-Shot Coordinate-Level Control in Complex Robotics Tasks: Our system operates without any prior training, fine-tuning, or environment-specific prompts, successfully handling diverse tasks in both simulated and real-world environments.
@@
  • Developed a Multi-Agent VLLM Framework to Overcome Previous Limitations in High-Level Robotic Planning from Language Instructions: We have developed a novel multi-agent structure within a single VLLM, where specialized agents collaboratively handle various aspects of high level robotic planning, from sequential position generation to location estimation. By integrating perception and action, and employing a divide-and-conquer approach with reflection capabilities, we address the shortcomings of previous models, including issues with context-aware object identification, precise localization, and handling multiple instances of the same object.
  • Empirical Validation through Extensive Experiments and Ablation Studies: We validate our framework with comprehensive experiments in both simulation (VIMABench) and real-world settings, including spatial planning tasks from Trajectory Generators and a new set of semantic reasoning real-world robotics tasks. Our results show significant performance improvements over existing methods on problems related to physical task planning. Interestingly, these improvements are also present when comparing against planning methods that require training. We also conduct thorough ablation studies to examine the effects of different agents and configurations, highlighting the critical role of the multi-agent system in achieving optimal performance.
@@ 2 Motivating Examples
2.1 Can an LLM as a Planner with a Separate Vision Model Find Objects?
Not Always. There are limitations at both the planning and perception levels, which are both intimately ralted.
At the planning level, non-vision LLMs cannot generate meaningful plans for ambiguous prompts that rely on environmental context. For example, consider the task: “Rank the fruits from most expensive to cheapest.” Without visual input to identify the fruits and their prices, the LLM cannot accurately rank them, nor generate useful queries for the vision model.
At the perception level, vision models also have limitations in context-aware perception. A notable prior work is the Trajectory Generator (Kwon et al., 2024), which uses GPT as a text planner and LangSAM as the vision model. In this approach, GPT extracts the objects to segment from the task prompt and passes them to LangSAM for object identification and segmentation. As illustrated in Figure 2, LangSAM fails to correctly identify or segment all intended objects based on the prompt. While this example highlights several challenges inherent in using separate vision models for complex tasks, it does not capture the full scope of limitations, which are discussed in detail below:
1. Difficulty with Less Common and Non-Segmented Objects: LangSAM struggles to identify uncommon objects (e.g., robot grippers, box lids) and abstract regions that cannot be clearly segmented. When objects are less prominent in the scene or when boundaries are not well-defined, LangSAM fails to provide accurate identification or spatial understanding.
2. Misinterpretation of Spatial and Positional Instructions: LangSAM often misinterprets vague spatial instructions like “pick up the rightmost object” due to its lack of precise spatial reasoning. In multi-instance scenarios, positional references like “the middle can” are challenging because the model frequently miscounts objects, leading to incorrect identification.
3. Lack of Contextual Awareness and Differentiation: LangSAM lacks the contextual understanding necessary to distinguish between relevant objects for manipulation and other elements in the scene. For instance, it may mistakenly select parts of the robot arm itself, failing to identify the intended target due to a lack of contextual awareness.
Figure 2: Examples of LangSAM’s detection failures in simulated environments. The bolded text within the prompts represents the objects extracted by GPT and passed to LangSAM.
Can These Issues Be Fixed?
Not within the current framework. Even with enhanced reasoning and replanning, we are unable to fully address LangSAM’s limitations because the LLM lacks the capability to detect, notice, or correct errors originating from the seperate vision model.
@@ 2.1 Can an LLM as a Planner with a Separate Vision Model Find Objects?
2.2 Could simply replacing LangSAM with a VLLM resolve these issues?
The answer to this question is ‘no.’ While replacing LangSAM with a VLLM may improve context comprehension, the resulting system fails to match the precision that LangSAM already provides. And thus the substitution is largely a wash. There are a few reasons for this.
@@ 2.2 Could simply replacing LangSAM with a VLLM resolve these issues?
However, while experimenting with using VLLMs to generate high-level plans, we had an interesting observation: VLLMs often know they’re wrong. And they can often diagnose their own errors.
@@
Even more interesting than this first observation, we also noticed that VLLMs often know how to iteratively self-correct errors in their perception. Over several iterations, they can improve their estimation of an object’s position, moving closer to the correct target (see Figure 4). This process of an LLM iteratively refine its outputs based on feedback is known as reflection.
Figure 4: An VLLM improving its estimation of the grapes’ position over several iterations.
Thus, we arrive at our main thesis. While naively using a VLLM for high-level robotic planning is often insufficient, the effectiveness of VLLMs for high-level robotic planning can be improved substantially if we leverage their self-correction capabilities in a structured way.
@@ 3 Wonderful Team
Building on insights from the previous section, we propose a novel pipeline for high-level physical task planning in robotics that leverages specialized agents, each responsible for a distinct part of the reasoning process within a structured framework. By combining the strengths of Vision-Language Models (VLLMs) and breaking down complex tasks into manageable components, each agent can focus on a specific role, resulting in more precise and reliable high-level planning. As illustrated in Figure 6, our multi-agent framework defines the distinct roles of each agent, the flow of information from high-level tasks to low-level actions, and their collaborative efforts in executing tasks effectively.
@@
We discuss the role of each team member and the flow of information between them below. Simplified prompts and more details on the implementation and workflow for each team member can be found in Appendix B. Full prompts are available in the project’s codebase.
@@
Figure 5: Overview of the major components of the Wonderful Team. Each part of the pipeline receives different levels of input, with a unique scope and specialization within the project. The agents collaboratively handle tasks ranging from high-level planning and logical verification to precise spatial reasoning and memory management, ensuring robust and efficient execution.
@@
Each agent in our system is designed to address specific challenges in high level planning for physical manipulation tasks. For example, in Figure 6(b), when the robot is instructed to “put the banana into the box,” the initial plan generated by the Supervisor agent often overlooks obstacles like the box’s lid. This is where the Verification agent plays a critical role. Its reflection process involves reviewing the subgoal plan, checking for potential issues such as physical constraints or incomplete steps, and cross-referencing this plan with the current state of the environment. If an issue, like the lid blocking access to the box, is detected, the Verification agent raises this concern to the Supervisor. This early feedback allows the system to refine the plan before executing any action.
@@
The Grounding team then takes over to refine the coordinates for each target, ensuring precise and collision-free movements. The Mover and Checker agents collaborate through an iterative process of adjusting positional groundings. Figure 4 provides an example of the Grounding team in action. The separation of tasks into a multi-agent system proves advantageous, as it allows each agent to focus on its distinct responsibilities with varying levels of access to critical information. For a detailed discussion on the benefits of this multi-agent approach, refer to Appendix F.3.3.
@@
Are all parts of the Wonderful Team necessary? Ablation studies reveal that all components of the Wonderful Team are essential. Removing memory agents leads to failures, such as mistaking irrelevant objects for targets, while omitting grounding members results in inaccurate coordinates. A supervisor-only setup works for simple tasks but fails with complex ones, lacking precision and corrective processes. Appendix D provides detailed analysis, and Table 11 in the appendix shows the impact on success rates when specific agents are removed. Appendix B shows the information flow from one agent to the next within the team.
@@
(b) A symbolic example illustrating the framework in (a).
@@ 4 Related Work
Language Models Empowering Robotics: Prior work has leveraged natural language to enhance robotic learning and adaptation. Early approaches equipped agents with learned language embeddings, requiring large amounts of training data (Bing et al., 2023; Jiang et al., 2023). Others focused on connecting language instructions with low-level action primitives to solve long-horizon tasks (Hu et al., 2023; Huang et al., 2022b; Liang et al., 2023). While effective in specific contexts, these methods often struggle to generalize to new tasks without retraining. Foundation VLA models like RT-1 (Brohan et al., 2022), RT-2 (Brohan et al., 2023), and OpenVLA (Kim et al., 2024) have advanced versatile robotic systems, but they still require significant training to achieve robust performance across diverse tasks. Additionally, from a system design perspective, directly mapping vision to action embeddings in these approaches is less intuitive compared to breaking down the problem into interpretable subtasks or subgoals, making it harder to intuitively reason about and improve the system’s behavior.
@@
Vision-Language Models for Localization: PIVOT (Nasiriany et al., 2024) addresses enabling VLMs to localize actionable points without fine-tuning on task-specific data. Their approach centers on localization through visual question answering, with minimal focus on planning—similar to the role of our Grounding Team. Unlike our method, which integrates both localization and planning within a multi-agent framework, PIVOT primarily addresses localization without managing complex, long-horizon tasks. In PIVOT, a single agent iteratively selects action points, whereas our approach employs multiple agents with distinct roles for refining and verifying actions. A detailed comparison is provided in Appendix F.2.
Language Models as Zero-Shot Trajectory Generators and MOKA: Recent works by Kwon et al. (2024) and Fang et al. (2024) demonstrate the use of pre-trained large language models (LLMs) for generating affordance-based, point-level subgoal trajectories. Both approaches rely on dedicated object detection systems that use a combination of GroundingDINO and SAM to extract object information, which informs LLM-based planning. These methods employ LLMs to generate executable Python scripts that define the robot’s trajectory. In contrast, our approach fundamentally differs by leveraging the same vision large language model for perception to seamlessly integrate planning and grounding without depending on external modules. We provide detailed comparisons in Appendices F.3 and F.4.
Natural Language as Policies: Concurrent with our work, Natural Language as Policies (NLaP) (Mikami et al., 2024) developed a few-shot, end-to-end model for coordinate-level action prediction. Their approach involves providing a one-shot example, either from the same task or a closely related one, rather than adopting a zero-shot paradigm. Unlike our method, which integrates both grounding and planning within a multi-agent framework, NLaP focuses less on grounding and directly uses system information from the environment, bypassing the need to extract coordinates from images using VLMs. NLaP serves as one of the baselines in our experiments, and a detailed comparison is presented in Appendix F.1.

### v6

@@ Abstract
We introduce Wonderful Team, a multi-agent Vision Large Language Model (VLLM) framework for executing high-level robotic planning in a zero-shot regime. In our context, zero-shot high-level planning means that for a novel environment, we provide a VLLM with an image of the robot’s surroundings and a task description, and the VLLM outputs the sequence of actions necessary for the robot to complete the task. Unlike previous methods for high-level visual planning for robotic manipulation, our method uses VLLMs for the entire planning process, enabling a more tightly integrated loop between perception, control, and planning. As a result, Wonderful Team’s performance on real-world semantic and physical planning tasks often exceeds methods that rely on separate vision systems. For example, we see an average 40% success rate improvement on VimaBench over prior methods such as NLaP, an average 30% improvement over Trajectory Generators on tasks from the Trajectory Generator paper, including drawing and wiping a plate, and an average 70% improvement over Trajectory Generators on a new set of semantic reasoning tasks including environment rearrangement with implicit linguistic constraints. We hope these results highlight the rapid improvements of VLLMs in the past year, and motivate the community to consider VLLMs as an option for some high-level robotic planning problems in the future.
@@ 1 Introduction
High-level robotic planning is a difficult task to pin down. It requires an extensive knowledge of the physical world. If planning is to be done via language instructions, it also requires a semantic understanding of the language itself, and how the objects in the environment relate to sometimes vague linguistic instructions. Consider, for example, the task “Pick up the object to the right of the grapes and put it in the box.” To solve this task, we need to identify ‘the object to the right of the grapes,’ which is a non-trivial visual relationship inference problem. We also need to consider the implicit constraints of the environment. For example, if the box has a closed lid, the task necessitates reasoning about lid removal before the object can be placed inside.
@@
Figure 1: Comparison of plans for a color-matching fruit placement task. MOKA’s trajectory (b) shows limitations arising from both unverified plans (e.g., step 2 treating the grape as green) and misalignment between visual grounding and VLLM planning (e.g., step 3 misidentifying the orange fruit as the ‘orange area’ and step 4 failing to pick the green apple due to unspecified color). These errors are often irrecoverable, even with VLLM correction (see Figure 15). See Appendix G.4 for details on the comparison with MOKA.
@@
Recently, Large Language Models (LLMs) have emerged as an enticing pathway for achieving high-level robotic planning. LLMs already have a built-in world model, including priors over object names and relative relationships. They also have the ability to understand and parse semantic meaning from language instructions. Several promising pipelines exist for converting language commands to high-level robotics plans Kwon et al. (2024). Typically, if vision-based planning is required, these pipelines make use of an off-the-shelf vision encoder such as LangSAM or OWL-ViT, which finds objects in an environment and delivers their coordinates in pixel space. While this is a great idea in theory, we find that, in practice, such reliance on an external vision system often results in pipelines that are brittle and unable to recover from mistakes originating in the visual system.
@@
In this paper, we consider the problem of using Visual Large Language Models (VLLMs) to generate action-level robotics plans directly from environmental images. Compared to prior works, we find that having VLLMs as the backbone of both planning and grounding in the pipeline offers several advantages. First, VLLMs combine the semantic reasoning power of language models with robust visual understanding, bridging the traditional gap between high-level planning and visual grounding as shown in Figure 1. More importantly, VLLMs enable self-correction at both the planning and perception levels. For planning, if the initial high-level plan overlooks environmental constraints, VLLMs can identify these errors and suggest revisions. At the perception level, while specialized detection models fail irreparably when objects cannot be found, our VLLM-based perception agents can refine their output through multimodal self-correction. By using updated visual feedback and annotations that include richer multimodal information, like the spatial relations between bounding boxes and nearby objects, the system improves its predictions and fixes errors more effectively. Our framework enhances error recovery by combining self-feedback with external guidance from higher-level agents when self-correction alone fails. This approach significantly improves task success rates, with real-world spatial and semantic planning tasks showing an approximate 80% increase compared to a baseline GPT-4o model without self-correction. Many tasks improve from 0% to 100% success rates with the introduction of self-correction. For detailed ablation studies on the impact of each component in the Wonderful Team framework, see Appendix E.
Our contributions can be summarized as follows:
@@
  • Zero-Shot Coordinate-Level Control in Complex Robotics Tasks: Our system operates without prior training, fine-tuning, or detailed task prompts and performs robustly across diverse domains while requiring minimal domain knowledge.
@@
  • Development of a Multi-Agent VLLM Framework to Overcome Limitations in Robotic Planning: We have developed a novel hierarchical multi-agent structure where specialized agents collaboratively handle various aspects of demanding robotics tasks, from high-level robotic planning to low-level position-based controlling. By integrating perception and action, enabling inter-agent communication, and employing a divide-and-conquer approach with reflection capabilities, we address the shortcomings of previous models, including challenges in long-horizon planning, implicit goal reasoning, context-aware object identification, irrecoverable perception failure, precise localization, and managing multiple instances of the same object.
  • Empirical Validation through Extensive Experiments and Ablation Studies: We validated our framework through extensive tabletop experiments, including long-horizon pick-and-place, sweeping, and fine-grained trajectory planning and generation, demonstrating its effectiveness beyond standard pick-and-place tasks. It outperformed other zero-shot approaches and fine-tuned vision-language action models, attributed to its self-correction mechanism and multi-agent system design, which enable robust performance in tasks requiring complex planning and precise object grounding. We additionally explored tasks outside our scope, including navigation and grounding in cluttered environments, showing our system’s potential for generalization. Detailed results and discussions are provided in Appendix B.
@@ 2 Motivating Examples
2.1 Can an LLM Planner with a Separate Vision Model Find Objects?
Not Always. Specialized detection models have limitations in context-aware perception. In both Trajectory Generator and MOKA, GPT extracts the objects to segment based on the task prompt and passes them to a separate detection model (LangSAM and GroundDINO, respectively) for object identification and segmentation. As illustrated in Figure 2, LangSAM and GroundingDINO both fail to correctly identify or segment target objects based on the provided prompt. While these examples highlight several challenges inherent in using separate vision models for complex tasks, they do not capture the full scope of limitations, which are discussed in detail below:
1. Difficulty with Less Common or Hard-to-Bound Objects: they struggle with identifying uncommon objects (e.g., robot grippers, box lids), as seen in Figure 2 (a), and abstract regions that cannot be clearly segmented. They also fail when objects are less prominent or boundaries are ill-defined.
2. Misinterpretation of Spatial and Positional Instructions: they often misinterpret vague spatial instructions like “pick up the rightmost object” due to its lack of precise spatial reasoning. In multi-instance scenarios, positional references like “the middle can” are challenging because the model frequently miscounts objects, leading to incorrect identification.
3. Lack of Contextual Awareness and Differentiation: they lack contextual understanding required to differentiate between similar targets based on instructions. For example, in Figure 1, when tasked with “place fruits in the corresponding colored region,” the model identifies the orange itself as both the “orange” and the “orange-colored region.” While this interpretation aligns with the literal prompt, it ignores the contextual requirement and results in execution failure.
Figure 2: Examples of detection failures by specialized segmentation models. Segmentation errors are common across prior methods, regardless of how the model is prompted. See Figure 15 for a detailed discussion on the extent of this issue and Appendix G for extensive comparisons with TG and MOKA.
Can These Issues Be Fixed Easily?
Even though Visual LLMs can verify results from specialized detection models, correcting detection errors remains a significant challenge. Separate detection models retain full control over object identification, limiting Visual LLMs to refining prompts with alternative object descriptions. As shown in Figure 2 (b) (and more extensively in Figure 15) larger changes—such as shifting from an object name (e.g., “soy sauce”) to a description of its shape or color (e.g., “tall, dark brown bottle”)—can sometimes improve results. However, the finite set of descriptors for any given object can lead to a dead-end. In some cases, this iterative process may become an inefficient loop, failing to isolate the desired object, especially when the optimal descriptor does not succeed in the given context.
@@ 2.1 Can an LLM Planner with a Separate Vision Model Find Objects?
2.2 Could simply replacing LangSAM with a VLLM solve these issues?
The answer to this question is ‘no.’ While replacing LangSAM with a VLLM may improve context comprehension, the resulting system fails to match the precision that LangSAM already provides. Thus, the substitution is largely ineffective. There are a few reasons for this.
@@ 2.2 Could simply replacing LangSAM with a VLLM solve these issues?
However, while experimenting with using VLLMs to generate high-level plans, we made an interesting observation: VLLMs often know they’re wrong, and they can often diagnose their own errors.
@@
Even more interestingly, we also noticed that VLLMs often know how to iteratively self-correct their perception errors. Over several iterations, they can improve their estimation of an object’s position, moving closer to the correct target (see Figure 4). This process of an LLM iteratively refining its outputs based on feedback is known as reflection.
Figure 4: A VLLM improving its estimation of the grapes’ position over several iterations.
Thus, we arrive at our main thesis. While naively using a VLLM for high-level robotic planning is often insufficient, leveraging their self-correction capabilities in a structured way can substantially improve their effectiveness.
@@ 3 Wonderful Team
Building on insights from the previous section, we propose a novel pipeline for high-level physical task planning in robotics that leverages specialized agents, each responsible for a distinct part of the reasoning process within a structured framework. By combining the strengths of Vision-Language Models (VLLMs) and breaking down complex tasks into manageable components, each agent can focus on a specific role, which results in more precise and reliable high-level planning. As illustrated in Figure 6, our multi-agent framework defines the distinct roles of each agent, the flow of information from high-level tasks to low-level actions, and their collaborative efforts in executing tasks effectively.
@@
We discuss the role of each team member and the flow of information between them below. Simplified prompts and more details on the implementation and workflow for each team member can be found in Appendix C. Full prompts are available in the project’s codebase.
@@
Figure 5: Overview of the major components of the Wonderful Team. Each part of the pipeline receives a different level of input, with a unique scope and specialization within the project. The agents collaboratively handle tasks ranging from high-level planning and logical verification to precise spatial reasoning and memory management, ensuring robust and efficient execution.
@@
Each agent in our system is designed to address specific challenges in high-level planning for physical manipulation tasks. For example, in Figure 6(b), when given the instruction ’put the banana into the box,’ the Supervisor agent’s initial plan often overlooks obstacles like the box’s lid. This is where the Verification agent plays a critical role. Its reflection process involves reviewing the subgoal plan, checking for potential issues, such as physical constraints or incomplete steps, and cross-referencing this plan with the current state of the environment. If an issue, like the lid blocking access to the box, is detected, the Verification agent raises this concern to the Supervisor. This early feedback allows the system to refine the plan before executing any action.
@@
The Grounding team then takes over to refine the coordinates for each target, ensuring precise and collision-free movements. The Mover and Checker agents collaborate through an iterative process of refining positional groundings. Figure 4 provides an example of the Grounding team in action. The separation of tasks into a multi-agent system proves advantageous, as it allows each agent to focus on its distinct responsibilities with varying levels of access to critical information. For a detailed discussion on the benefits of this multi-agent approach, refer to Appendix G.3.3.
@@
Are all components of the Wonderful Team necessary? Ablation studies show that all components of the Wonderful Team are essential. Removing memory agents leads to failures, such as mistaking irrelevant objects for targets, while omitting grounding agents results in inaccurate coordinates. A supervisor-only setup works for simple tasks but fails with complex ones, lacking precision and self-correction mechanisms. Appendix E provides detailed analysis, and Table 11 shows success rate impacts. Appendix C illustrates the team’s information flow.
@@
(b) A symbolic example illustrating the framework in Figure 6(a).
@@ 4 Related Work
Language Models Empowering Robotics: Prior work has leveraged natural language to enhance robotic learning and adaptation. Early approaches equipped agents with learned language embeddings, requiring large amounts of training data (Bing et al., 2023; Jiang et al., 2023). Others focused on connecting language instructions with low-level action primitives to solve long-horizon tasks (Hu et al., 2023; Huang et al., 2022b; Liang et al., 2023). While effective in specific contexts, these methods often struggle to generalize to new tasks without retraining. Foundation VLA models like RT-1 (Brohan et al., 2022), RT-2 (Brohan et al., 2023), and OpenVLA (Kim et al., 2024) have advanced versatile robotic systems, but they still require significant training to achieve robust performance across diverse tasks. Additionally, from a system design perspective, directly mapping vision to action embeddings in these approaches is less intuitive compared to breaking down the problem into interpretable subtasks or subgoals, making it harder to intuitively reason about and improve the system’s behavior. For a detailed comparison with RT-1 and OpenVLA, see Appendix G.5.
@@
Vision-Language Models for Localization: PIVOT (Nasiriany et al., 2024) enables VLMs to localize actionable points without fine-tuning on task-specific data. Their approach centers on localization through visual question answering, with minimal focus on planning—similar to the role of our Grounding Team. Unlike our method, which integrates both localization and planning within a multi-agent framework, PIVOT primarily addresses localization without managing complex, long-horizon tasks. In PIVOT, a single agent iteratively selects action points, whereas our approach employs multiple agents with distinct roles for refining and verifying actions. A detailed comparison is provided in Appendix G.2.
Language Models as Zero-Shot Trajectory Generators and MOKA: Recent works by Kwon et al. (2024) and Fang et al. (2024) demonstrate the use of pre-trained large language models (LLMs) for generating affordance-based, point-level subgoal trajectories. Both approaches rely on dedicated object detection systems that use a combination of GroundingDINO (Liu et al., 2023b) and SAM (Kirillov et al., 2023) to extract object information, which informs LLM-based planning. These methods employ LLMs to generate executable Python scripts that define the robot’s trajectory. In contrast, our approach fundamentally differs by leveraging the same VLLM for perception to seamlessly integrate planning and grounding without depending on external modules. We provide detailed comparisons in Appendices G.3 and G.4.
Natural Language as Policies: Concurrent with our work, Natural Language as Policies (NLaP) (Mikami et al., 2024) developed a few-shot, end-to-end model for coordinate-level action prediction. Their approach involves providing a one-shot example, either from the same task or a closely related one, rather than adopting a zero-shot paradigm. Unlike our method, which integrates both grounding and planning within a multi-agent framework, NLaP focuses less on grounding and directly uses system information from the environment, bypassing the need to extract coordinates from images using VLMs. NLaP serves as one of the baselines in our experiments, and a detailed comparison is presented in Appendix G.1.